### v1

# Temporal Entanglement Transitions in the Periodically Driven Ising Chain

Abstract
Periodically driven quantum systems can host non-equilibrium phenomena without static analogs, including in their entanglement dynamics. Here, we discover temporal entanglement transitions in a Floquet spin chain, which correspond to a quantum phase transition in the spectrum of the entanglement Hamiltonian and are signaled by dynamical spontaneous symmetry breaking. We show that these transitions are entanglement-driven, i.e., they require initially entangled states and remain invisible to conventional local observables. Intriguingly, we find these transitions across a broad range of driving frequencies (from adiabatic to high-frequency regime) and independently of drive details, where they manifest as periodic, sharp entanglement spectrum reorganizations marked by the Schmidt-gap closure, a vanishing entanglement echo, and symmetry-quantum-number flips. At high frequencies, the entanglement Hamiltonian acquires an intrinsic timescale decoupled from the drive period, rendering the transitions genuine steady-state features. Finite-size scaling reveals universal critical behavior with correlation-length exponent $\nu=1$, matching equilibrium Ising universality despite its emergence from purely dynamical mechanisms decoupled from static criticality. Our work establishes temporal entanglement transitions as novel features in Floquet quantum matter.
Introduction.— Periodically driven quantum many-body systems are a powerful platform for exploring quantum phases inaccessible in equilibrium, opening the door to the coherent control and engineering of quantum matter via time-periodic fields [1, 2, 3, 4, 5]. The study of Floquet driven systems has unveiled remarkable phenomena ranging from Floquet topological insulators [6, 7] and time crystals [8, 9, 10] to dynamical localization [11, 12] and prethermalization [13, 14, 15]. While extensive research has focused on conventional observables like magnetization and transport properties, exploring the entanglement structure of driven quantum systems remains an active area of research, given its fundamental role in characterizing quantum phases and phase transitions [16, 17]. Of particular interest is the entanglement spectrum (ES) [18], which provides direct access to the entanglement Hamiltonian and encodes information beyond that contained in the entanglement entropy alone. The ES serves as a powerful diagnostic for capturing the universal features of quantum phases in equilibrium, especially in the context of gapped ground states but also for critical points [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. There is growing recognition that the ES also reveals distinctive signatures of quantum chaos, thermalization, and criticality out of equilibrium [41, 42, 43, 44].
Recent work has begun exploring the interplay between driving and entanglement, including for Floquet-driven conformal field theories [45, 46, 47] and in driven-dissipative systems [48, 49, 50]. Particularly relevant are studies of Page curves in unitary dynamics [51, 52, 53], where entanglement transitions emerge from the competition between entanglement generation and relaxation processes in the presence of conserved charges. While equilibrium quantum phase transitions are intimately connected to entanglement scaling and the structure of the many-body ground state [54, 55, 17], the analogous relationship in driven systems remains unexplored. Since driven systems can access dynamical phases with no equilibrium counterpart, this suggests new entanglement phenomena are waiting to be uncovered. Indeed, developments in geometric Floquet theory [56, 57], advances in conformal field theory approaches to quantum quenches [58, 59, 60], and emerging concepts of entanglement transitions [61, 62, 63, 64, 65, 66, 67] have unearthed novel far-from-equilibrium entanglement dynamics and, in many cases, connections with information-theoretic quantities [68, 69, 70, 71]. However, whether non-analytic entanglement transitions can occur in generic periodically-driven systems remains an open question.
In this Letter, we report the discovery of temporal entanglement transitions via dynamical spontaneous symmetry breaking in the entanglement Hamiltonian (EH) of a periodically driven spin chain. These transitions are uniquely entanglement-driven, manifesting only for initially entangled states and eluding conventional probes such as magnetization or Loschmidt echoes. They manifest as (1) periodic, sharp reorganizations of the entanglement spectrum marked by Schmidt gap closure [72, 73, 74], (2) vanishing overlap between initial and instantaneous entanglement ground states (also known as the entanglement echo [75]), and (3) symmetry quantum number flips [76, 77, 78]. Crucially, these transitions occur across all driving frequencies and exhibit universal critical behavior with correlation length exponent $\nu=1$, matching the equilibrium Ising universality class [79, 80], despite emerging from purely dynamical mechanisms and decoupled from any underlying equilibrium criticality.
Our findings establish temporal entanglement transitions as a distinct class of non-equilibrium phenomenon that is uniquely detectable through entanglement measures while remaining hidden from conventional observables, revealing that the EH itself can undergo quantum phase transitions. At high frequencies, we find that the EH gains an intrinsic timescale decoupled from the drive, rendering its Floquet-periodic transitions as genuine steady-state features. The universal critical behavior persists across different equilibrium phases of the initial state, confirming the transitions’ fundamentally non-equilibrium character, decoupled from the underlying equilibrium phase diagram.
Figure 1: Temporal entanglement transitions in the driven TFIM. (a) Schmidt gap $\Delta\lambda=\lambda_{0}-\lambda_{1}$ closing periodically at critical times. (b) Entanglement echo $|E(t)|^{2}$ vanishing at odd critical times $t_{c}^{(k=\text{odd})}$, as the dominant Schmidt vector $|\lambda_{0}\rangle$ orthogonalizes into a distinct symmetry sector. (c) Subsystem parity expectations showing dynamical spontaneous $\mathbb{Z}_{2}$ symmetry breaking in the entanglement ground state (green) and complementary behavior in the first-excited state (orange). The alternating parity pattern provides consistent physical explanation of the echo’s behavior. Red vertical lines mark critical times of symmetry breaking, and overlaid across all panels to provide a synchronized, consistent picture. Parameters: $L=24$, $L_{A}=9$, $J=1.0$, $h_{0}=2.0$, $\omega=5.0$, $dt=0.01$.
Model and Results.— We consider the transverse-field Ising model (TFIM) with open boundary conditions, consisting of $L$ spin-$1/2$ particles with time-dependent Hamiltonian
${\cal H}(t)=-J\sum_{i=1}^{L-1}\sigma_{i}^{z}\sigma_{i+1}^{z}-h(t)\sum_{i=1}^{L}\sigma_{i}^{x},$ (1)
where $J$ is the Ising coupling, $h(t)=(h_{0}/2)\cos(\omega t)$ is the oscillating transverse field, and $\sigma_{i}^{\alpha}$ are Pauli matrices. We initialize the system in the ground state of the static Hamiltonian ${\cal H}_{\text{static}}={\cal H}(t)|_{t=0}$ and evolve under ${\cal H}(t)$ using time-dependent variational principle (TDVP) [81, 82, 83] for matrix product states (MPS) [84], as implemented in the ITensors library [85] (see Supplemental Material (SM) for details [86]). Note that our results are independent of the specific choice of Floquet drive.
Our primary focus is the entanglement dynamics of subsystem $A$, comprising the first $L_{A}$ spins. From the reduced density matrix $\rho_{A}=\text{Tr}_{\bar{A}}|\psi(t)\rangle\langle\psi(t)|$ ($\bar{A}$ is the remaining $L-L_{A}$ sites), we extract Schmidt values $\lambda_{i}(t)$} (which satisfies $\sum_{i}\lambda_{i}=1$ at each instant of time $t$) and define the entanglement Hamiltonian (EH) ${\cal H}_{\text{ent}}(t)$ via $\rho_{A}(t)\equiv e^{-{\cal H}_{\text{ent}}(t)}$ [18, 39]. Since they commute, the eigenstates are shared by ${\cal H}_{\text{ent}}(t)$ and $\rho_{A}(t)$ at each instant of time. We probe the dynamics of ${\cal H}_{\text{ent}}(t)$ by examining its instantaneous eigenspectrum, ${\cal H}_{\text{ent}}(t)|\lambda_{n}(t)\rangle=\epsilon_{n}(t)|\lambda_{n}(t)\rangle$, and the expectation values of subsystem operators within these states. Here, the index $n=0,1,2,\ldots$ labels the eigenstates with eigenvalues $\epsilon_{n}$ in ascending order, where $\epsilon_{0}$ denotes the “ground-state” eigenvalue. The transformation $\lambda_{n}=e^{-\epsilon_{n}}$ gives the eigenvalues of the reduced density matrix $\rho_{A}$ (the Schmidt values), with $\lambda_{0}$ being the largest. Note that the reduced density matrix $\rho_{A}$ allows us to evaluate all Rényi entropies $S_{n}\left(\rho_{A}\right)=\frac{1}{1-n}\ln\left(\sum_{i=1}^{r}\lambda_{i}^{n}\right)$, where the min-entropy ($n\to\infty$) is given by $S_{\text{min}}=-\ln\lambda_{0}=\epsilon_{0}$ (ground state energy of the EH). Moreover, a fictitious temperature $1/n$ can be associated to the EH (see Ref. [52]).
A crossing in the two largest Schmidt values, $\lambda_{0}(t)$ and $\lambda_{1}(t)$, signifies a fundamental reorganization of the ES, corresponding to a non-analyticity in the ground state energy $\epsilon_{0}(t)$ of ${\cal H}_{\text{ent}}(t)$ (with fictitious temperature $T_{\text{fict}}=\lim_{n\to\infty}\frac{1}{n}=0$), signaling a temporal quantum phase transition in the EH. Surprisingly, we observe these transitions across all non-zero driving frequencies (from $\omega\ll$ $\min(h_{0},J)$ to $\omega\gg$ $\max(h_{0},J)$), indicating the generic nature of this phenomenon in the periodically-driven TFIM across all frequencies. This behavior emerges specifically when the initial state is entangled (such as the ground state of ${\cal H}_{\text{static}}$), rather than a simple product state or a domain wall state with vanishing entanglement at the subsystem boundary (see SM [86]).
To diagnose dynamical transitions within the entanglement spectrum, we introduce and track specific observables designed to detect symmetry breaking. The central quantity is the entanglement echo $E(t)=\langle\lambda_{0}(0)|\lambda_{0}(t)\rangle$, which measures the fidelity of the instantaneous entanglement ground state to its initial configuration [75]. A vanishing echo signals an orthogonalization of $|\lambda_{0}(t)\rangle$, suggesting a crossing into a distinct symmetry sector. To directly test this, we compute the expectation values of symmetry operators within the entanglement eigenstates. Specifically, we monitor the subsystem parity $\langle\lambda_{n}(t)|P_{A}|\lambda_{n}(t)\rangle$, where $P_{A}=\prod_{i\in A}\sigma_{i}^{x}$. A spontaneous change in the parity of the dominant state $|\lambda_{0}(t)\rangle$, coinciding with a vanishing entanglement echo and a vanishing Schmidt gap $\lambda_{0}-\lambda_{1}$, constitutes the hallmark of a temporal entanglement transition. For each subsequent interval, the dynamics alternate between two dwell times: $T_{-}$, during which the dominant state carries $\langle P_{A}\rangle=-1$, and $T_{+}$, during which it carries $\langle P_{A}\rangle=+1$. At intermediate drives ($\omega\gtrsim$ 7), these intervals become individually regular yet remain unequal, $T_{-}\neq T_{+}$, evidencing partial “Floquet inheritance” by the EH. For higher frequencies ($\omega\gtrsim$ 10), the alternation persists but the intervals synchronize, $T_{-}\approx T_{+}\equiv T_{c}$, and $T_{c}$ saturates to an $\omega$-independent constant value (Fig. 3). This synchronization and saturation indicate complete Floquet inheritance: the EH dynamics are governed by a high-frequency effective (Floquet-Magnus) description, in which the entanglement transitions form a uniformly spaced temporal lattice set by an emergent, drive-induced time-scale rather than the bare period. See Table I and Sec. III.B in SM [86] for further details.
Unless otherwise noted, we use $L=24$ (open boundaries), $L_{A}=4$1212, $J=1.0$, $h_{0}=2.0$ (corresponding to equilibrium criticality), and $\omega=5.0$. Time steps $dt\leq 0.1/\omega$ ensure numerical stability; convergence is verified in SM [86].
Temporal Entanglement Transitions.— The combined signatures of a temporal entanglement transition are unambiguously observed in the dynamics of the driven TFIM. Fig. 1 presents a comprehensive view of the entanglement dynamics for a subsystem of size $L_{A}=9$, revealing periodic occurrences of the transition.
The onset of a transition is heralded by a critical slowdown of the entanglement spectrum, manifested as a sharp narrowing of the Schmidt gap $\Delta\lambda=\lambda_{0}-\lambda_{1}$ (Fig. 1a). The gap closes to within numerical precision at multiple critical times $t_{c}^{(k)}$}, where we denote the first critical time as $t_{c}^{(k=1)}\equiv t^{*}$, indicating a near-degeneracy of the two largest Schmidt values. This closing gap is the direct signature of an impending energy-level crossing in the EH ${\cal H}_{\text{ent}}(t)$ (recall that the largest and the second largest Schmidt values directly correspond to the ground state and first-excited state energies of ${\cal H}_{\text{ent}}(t)$).
Concurrently, the entanglement echo $E(t)=\left\langle\lambda_{0}(0)\mid\lambda_{0}(t)\right\rangle$ exhibits sharp vanishing at odd critical times (Fig. 1b). At these times $t_{c}^{(k=\text{odd})}$, where the Schmidt gap closes, $|E(t)|^{2}$ drops to zero, establishing that the instantaneous entanglement ground state $|\lambda_{0}(t_{c})\rangle$ orthogonalizes relative to its initial configuration. This confirms the reorganization of quantum correlations and the transition into a distinct symmetry sector of the Hilbert space.
The nature of this sector is revealed by the expectation value of the subsystem parity operator $P_{A}$ (Fig. 1c). For $t<t^{*}$, the ground state of ${\cal H}_{\text{ent}}(t)$ exhibits a well-defined parity of $\langle P_{A}\rangle=+1$. At each transition point $t^{*}$, this parity expectation flips discontinuously to $\langle P_{A}\rangle=-1$, indicating a dynamical spontaneous breaking of the $\mathbb{Z}_{2}$ symmetry by the entanglement ground state. Notably, the first excited state $|\lambda_{1}(t)\rangle$ displays the opposite parity, confirming that the gap closing corresponds to an exchange of roles between two symmetry-broken states. The subsystem magnetization $\langle M_{A}\rangle$ [80] and Loschmidt echo rate function [87, 88, 89, 90] (verified in SM [86]) remain smooth through these events, confirming the transition is of a purely entanglement-driven, dynamical spontaneous symmetry-breaking character.
We observe these transition signatures across all non-zero driving frequencies $\omega$ studied, indicating the generic nature of this phenomenon in periodically driven Ising chain. This behavior emerges specifically when the initial state is entangled (such as the ground state of ${\cal H}_{\text{static}}$), rather than a simple non-entangled state (see SM [86]). While transitions occur at all frequencies, the periodic regularity of critical times $t_{c}^{(k)}$} becomes pronounced only at higher driving frequencies, suggesting a crossover to Floquet-like behavior at the level of the EH (details provided in the SM [86]).
The synchronization of these three diagnostics, namely the closing of the Schmidt gap, the vanishing of the entanglement echo, and the discontinuous flip in parity, provides definitive evidence of a temporal entanglement transition. These are not independent events but are intrinsically linked manifestations of the same underlying phenomenon: a periodically-driven, dynamical quantum phase transition of the EH ${\cal H}_{\text{ent}}(t)$.
Finite-Size Scaling.— The temporal entanglement transitions are marked by non-analytic kinks in the ground state energy $\epsilon_{0}(t)=-\ln\lambda_{0}(t)$ of the EH. To establish the critical nature and universality of temporal entanglement transitions, we perform comprehensive finite-size scaling analysis across subsystem sizes $L_{A}=4,5,\ldots,12$.
Fig. 2 demonstrates the scaling properties of the first critical time $t_{c}^{(1)}=t^{*}$ and the minimum Schmidt gap density $S_{\text{min}}/L_{A}$ at criticality. The critical time exhibits power-law scaling $t^{*}/L_{A}\propto L_{A}^{-1/\nu}$ with critical exponent $\nu\simeq 1.00$ (Fig. 2a) that corresponds to the divergence of correlation length. Simultaneously, the critical entropy density follows $S_{\text{min}}/L_{A}\propto L_{A}^{-a}$ with $a\simeq 1$ (Fig. 2b). Surprisingly, the correlation length exponent $\nu=1$ establishes this as a continuous quantum phase transition for all ranges of driving frequencies, belonging to the same universality class as the equilibrium 2D classical Ising/1D quantum TFIM. Details provided in the SM [86] confirm that this exponent match is coincidental: to the best of our knowledge, temporal entanglement transitions appear to occur independently of the equilibrium phase diagram.
The universality of these transitions is demonstrated through data collapse using the scaling ansatz (recall $\epsilon_{0}=S_{\text{min}}$):
$\frac{\epsilon_{0}}{L_{A}}=\frac{1}{L_{A}^{a}}\mathcal{F}\left[\left(\frac{t}{L_{A}}-\frac{t_{c}}{L_{A}}\right)L_{A}^{1/\nu}\right],$ (2)
where $\mathcal{F}$ is a universal scaling function, and $\nu$ is the critical exponent corresponding to diverging correlation length. Fig. 2c shows excellent collapse of data from all subsystem sizes onto a universal curve, validating the scaling hypothesis. The agreement is further verified in Fig. 2d, which shows the raw data near the critical point.
Significantly, the same exponent describes all subsequent kinks at $t_{c}^{(k\geq 1)}$ for higher driving frequencies ($\omega\gtrsim 5.0$ for our parameter choices), where the EH inherits Floquet-like periodicity. This universality across multiple transitions suggests a common underlying fixed point. The exponent $\nu\simeq 1.00$ defines a new universality class for non-equilibrium entanglement dynamics in driven systems, which appears to be completely decoupled from the underlying equilibrium criticality (more on this below).
Figure 2: Finite-size scaling analysis of temporal entanglement transitions. (a) Critical time scaling $t^{*}/L_{A}\propto L_{A}^{-1/\nu}$ with $\nu=1.00$. (b) Critical entropy density scaling $\epsilon_{0}/L_{A}\propto L_{A}^{-a}$ with $a=1.00$. (c) Universal data collapse using scaling ansatz in Eq. (2). (d) Raw data before scaling collapse. The exponent $\nu=1$ establishes the same universality class for our driven non-equilibrium entanglement spectrum as equilibrium 2D classical Ising/1D TFIM. Parameters: $L=24$, $J=1.0$, $h_{0}=2.0$, $\omega=5.0$, $dt=0.01$.
Figure 3: Frequency dependence of the first critical time $t^{*}$ across subsystem sizes $L_{A}=4-12$. At low frequencies, $t^{*}\propto\omega^{-1}$ (adiabatic regime), while at high frequencies ($\omega\gtrsim 10$), $t^{*}$ saturates to frequency-independent values determined solely by subsystem size, indicating a crossover to Floquet steady-state behavior where the EH develops an intrinsic timescale. The universal saturation indicates that temporal entanglement transitions become intrinsic properties of the effective time-independent dynamics rather than driven phenomena, validating the Floquet-Magnus description and establishing characteristic timescales independent of the external driving protocol. Parameters: $L=24$, $J=1.0$, $h_{0}=2.0$. Time steps are chosen as $dt=0.01$ for $\omega=[0.1,10.0]$, $dt=0.002$ for $\omega=\{30,50\}$ and $dt=0.001$ for $\omega=\{70,100\}$. A total of 153 data points in this plot show excellent collapse (see Fig. 9 of SM [86]).
Effective Steady State at High Frequency.— At high driving frequencies, the temporal entanglement transitions can be understood through Floquet-Magnus theory [91, 92, 93]. We decompose the time-dependent Hamiltonian as $\mathcal{H}(t)=\mathcal{A}+\mathcal{B}\cos(\omega t)$, where $\mathcal{A}=-J\sum_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}$ and $\mathcal{B}=-(h_{0}/2)\sum_{i}\sigma_{i}^{x}$. The effective time-independent Hamiltonian emerges as a systematic expansion $\mathcal{H}_{\text{eff}}=\mathcal{H}_{0}+\mathcal{H}_{1}+\mathcal{H}_{2}+\cdots$ ${\cal H}_{(n)}$ is of order ${\cal O}(\omega^{-n})$], where the leading corrections are
${\cal H}_{\mathrm{eff}}=$ $-J\left(1+\frac{h_{0}^{2}}{2\omega^{2}}\right)\sum_{i=1}^{L-1}\sigma_{i}^{z}\sigma_{i+1}^{z}+\frac{h_{0}^{2}J}{2\omega^{2}}\sum_{i=1}^{L-1}\sigma_{i}^{y}\sigma_{i+1}^{y}$ (3)
$-\frac{2h_{0}J^{2}}{\omega^{2}}\left(\sigma_{1}^{x}+\sigma_{L}^{x}\right)-\frac{4h_{0}J^{2}}{\omega^{2}}\sum_{i=2}^{L-1}\sigma_{i}^{x}$
$-\frac{4h_{0}J^{2}}{\omega^{2}}\sum_{i=2}^{L-1}\sigma_{i-1}^{z}\sigma_{i}^{x}\sigma_{i+1}^{z}+{\cal O}(\omega^{-3}).$
Beyond renormalizing the Ising coupling, the expansion generates new terms: YY interactions, enhanced transverse fields at boundaries, and crucially, three-body interactions $\sim\sigma_{i-1}^{z}\sigma_{i}^{x}\sigma_{i+1}^{z}$ that couple neighboring bonds through local spin flips.
As demonstrated in SM [86], exact time evolution under $\mathcal{H}(t)$ shows good agreement with evolution under $\mathcal{H}_{\text{eff}}$ at higher frequencies, with entanglement entropy, Schmidt gap dynamics, and state fidelity exhibiting near-perfect overlap. Temporal entanglement transitions persist in the effective evolution with identical critical exponent $\nu$, confirming that these phenomena represent genuine features of the driven steady state where the EH develops an intrinsic timescale.
Table 1: Conditions for temporal entanglement transitions across different scenarios.
System
Initial state
(Product/Entangled)
Transitions
Equilibrium TFIM Any No
Driven TFIM Product No
Driven TFIM Entangleda Yes
Conserved‐charge
models
Any Yesb
a See SM for detailed discussion on dependence on initial state. b Only when $\rho_{A}$ is forced between disconnected sectors [51, 52].
Discussion and Conclusion.— We have demonstrated that the Floquet driven Ising chain hosts temporal entanglement transitions characterized by dynamical spontaneous symmetry breaking within the EH. These transitions exhibit universal critical behavior with exponent $\nu=1$, identical to equilibrium quantum criticality in the transverse-field Ising universality class, yet emerge from non-equilibrium dynamics that are decoupled from static criticality, as shown by their persistence across different equilibrium phases ($h_{0}=1.6$ to $2.4$, see SM [86]).
This work establishes that: (1) the ES develops non-analyticities for all $\omega$, with periodic reorganizations at higher frequencies; (2) transitions are diagnosed by synchronized Schmidt gap closure, entanglement echo vanishing, and parity flipping, evidencing dynamical spontaneous symmetry breaking; (3) finite-size scaling reveals universal $\nu=1$ critical behavior across frequencies; (4) at high frequencies, Floquet-Magnus theory provides an effective description accurately capturing the transition; (5) as summarized in Table 1, transitions require both driving and initial entanglement, distinguishing them from transitions found for initial product states [53] (more details in the SM [86]).
Fig. 3 shows critical times $t^{*}$ saturating to $\omega$-independent values at high frequencies, indicating intrinsic timescales of the driven steady state. The collapse of $t^{*}$ curves across subsystem sizes demonstrates universal Floquet inheritance, where the EH acquires periodicity and full data collapse is satisfied by the same $\nu=1$ exponent for all transitions [86].
These findings establish temporal entanglement transitions as fundamental aspects of non-equilibrium Floquet quantum dynamics. Notably, previous work on Floquet symmetry-protected topological system has demonstrated band crossings in entanglement spectrum [94], however the question of robust universality has been an open question. This is what we have addressed in this work, where our temporal entanglement transitions arise universally for any initially entangled state (topological or trivial) and exhibit nonanalytic reorganization (quantum phase transition) of the EH with Ising‐class critical exponents $\nu=1$. The observed behavior suggests driven systems can host rich critical phenomena as diagnosed through their entanglement dynamics. Temporal entanglement transitions define a class of non-equilibrium criticality distinct from Loschmidt echo singularities and magnetization dynamics (see SM [86]). The exact Jordan-Wigner equivalence between the TFIM and Kitaev chain suggests that these transitions should also manifest in driven topological superconductors, opening pathways to entanglement-based probes of Floquet Majorana physics [95, 96]. The requirement for initial entangled states motivates future exploration of how entanglement resources control critical structures in quantum resource theories [97]. Future directions include extensions to higher dimensions, long-range interactions, and potential realization in cold-atoms or trapped ion experiments [98, 99, 100].
Acknowledgments.— K. G. and R. J. acknowledge financial support by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Grants No. 436382789, and No. 493420525, via large equipment grants (GOEGrid). This work is supported by the Office of Advanced Scientific Computing Research, Exploratory Research for Extreme Scale Science (EXPRESS) program, Office of Science of the U.S. Department of Energy under Award Number DE-SC0026216 (A. P.). This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). A. P. thanks the Kavli Institute for Theoretical Physics (KITP) for its hospitality during the program on “Noise-robust Phases of Quantum Matter,” during which part of this work was completed.
Data and code availability.— All data and code used for data generation are available on Zenodo on reasonable request [101].

### v2

@@
 Periodically driven quantum systems can host non-equilibrium phenomena without static analogs, including in their entanglement dynamics.
-Here, we discover temporal entanglement transitions in a Floquet spin chain, which correspond to a quantum phase transition in the spectrum of the entanglement Hamiltonian and are signaled by dynamical spontaneous symmetry breaking.
-We show that these transitions are entanglement-driven, i.e., they require initially entangled states and remain invisible to conventional local observables.
-Intriguingly, we find these transitions across a broad range of driving frequencies (from adiabatic to high-frequency regime) and independently of drive details, where they manifest as periodic, sharp entanglement spectrum reorganizations marked by the Schmidt-gap closure, a vanishing entanglement echo, and symmetry-quantum-number flips.
+Here, we discover temporal entanglement transitions (TET) in a Floquet spin chain, which correspond to a quantum phase transition in the spectrum of the entanglement Hamiltonian and are signaled by dynamical spontaneous symmetry breaking.
+We identify the symmetry principles underlying these transitions: they appear when the driven Hamiltonian preserves global symmetry (here, $\mathbb{Z}_{2}$), the initial state respects this symmetry, and the reduced density matrix carries weight in both subsystem-parity sectors, with TET occurring precisely when the sector weights become equal (given the previous two conditions are also satisfied).
+Intriguingly, we find these transitions across a broad range of driving frequencies (from adiabatic to high-frequency regime) and independently of drive details, where they manifest as periodic, sharp entanglement spectrum reorganizations marked by the Schmidt-gap closure, a vanishing entanglement echo, and symmetry-quantum-number flips, while remaining invisible to conventional local observables.
 At high frequencies, the entanglement Hamiltonian acquires an intrinsic timescale decoupled from the drive period, rendering the transitions genuine steady-state features.
 Finite-size scaling reveals universal critical behavior with correlation-length exponent $\nu=1$, matching equilibrium Ising universality despite its emergence from purely dynamical mechanisms decoupled from static criticality.
-Our work establishes temporal entanglement transitions as novel features in Floquet quantum matter.
+Our work establishes TET as novel features in Floquet quantum matter.
 Introduction.— Periodically driven quantum many-body systems are a powerful platform for exploring quantum phases inaccessible in equilibrium, opening the door to the coherent control and engineering of quantum matter via time-periodic fields [1, 2, 3, 4, 5].
@@
 There is growing recognition that the ES also reveals distinctive signatures of quantum chaos, thermalization, and criticality out of equilibrium [41, 42, 43, 44].
-Recent work has begun exploring the interplay between driving and entanglement, including for Floquet-driven conformal field theories [45, 46, 47] and in driven-dissipative systems [48, 49, 50].
-Particularly relevant are studies of Page curves in unitary dynamics [51, 52, 53], where entanglement transitions emerge from the competition between entanglement generation and relaxation processes in the presence of conserved charges.
+Recent work has begun exploring the interplay between driving and entanglement, including for Floquet-driven conformal field theories [45, 46, 47, 48] and in driven-dissipative systems [49, 50, 51].
+Particularly relevant are studies of Page curves [52, 53, 54], where entanglement transitions emerge from competition between entanglement generation and relaxation in the presence of conserved charges.
 While equilibrium quantum phase transitions are intimately connected to entanglement scaling and the structure of the many-body ground state [55, 56, 17], the analogous relationship in driven systems remains unexplored.
-Since driven systems can access dynamical phases with no equilibrium counterpart, this suggests new entanglement phenomena are waiting to be uncovered.
-Indeed, developments in geometric Floquet theory [56, 57], advances in conformal field theory approaches to quantum quenches [58, 59, 60], and emerging concepts of entanglement transitions [61, 62, 63, 64, 65, 66, 67] have unearthed novel far-from-equilibrium entanglement dynamics and, in many cases, connections with information-theoretic quantities [68, 69, 70, 71].
+Since driven systems can access dynamical phases with no equilibrium counterpart, this suggests new entanglement phenomena waiting to be uncovered.
 However, whether non-analytic entanglement transitions can occur in generic periodically-driven systems remains an open question.
-In this Letter, we report the discovery of temporal entanglement transitions via dynamical spontaneous symmetry breaking in the entanglement Hamiltonian (EH) of a periodically driven spin chain.
-These transitions are uniquely entanglement-driven, manifesting only for initially entangled states and eluding conventional probes such as magnetization or Loschmidt echoes.
-They manifest as (1) periodic, sharp reorganizations of the entanglement spectrum marked by Schmidt gap closure [72, 73, 74], (2) vanishing overlap between initial and instantaneous entanglement ground states (also known as the entanglement echo [75]), and (3) symmetry quantum number flips [76, 77, 78].
-Crucially, these transitions occur across all driving frequencies and exhibit universal critical behavior with correlation length exponent $\nu=1$, matching the equilibrium Ising universality class [79, 80], despite emerging from purely dynamical mechanisms and decoupled from any underlying equilibrium criticality.
-Our findings establish temporal entanglement transitions as a distinct class of non-equilibrium phenomenon that is uniquely detectable through entanglement measures while remaining hidden from conventional observables, revealing that the EH itself can undergo quantum phase transitions.
+In this Letter, we report the discovery of temporal entanglement transitions (TET) via dynamical spontaneous symmetry breaking in the entanglement Hamiltonian (EH) of a periodically driven spin chain.
+These transitions require $\mathbb{Z}_{2}$-symmetric driven Hamiltonian and initial state, with critical times occurring at sector-weight degeneracies (made precise below), and manifest through three synchronized signatures: Schmidt gap closure [57, 58, 59], vanishing entanglement echo [60], and parity flips [61, 62, 63], while remaining invisible to conventional observables (magnetization, Loschmidt echo).
+Crucially, these transitions occur across all driving frequencies and exhibit universal critical behavior with correlation length exponent $\nu=1$, matching the equilibrium Ising universality class [64, 65], despite emerging purely dynamically, decoupled from any underlying equilibrium criticality.
+Our findings establish TET as a distinct class of non-equilibrium phenomenon that is uniquely detectable through entanglement measures while remaining hidden from conventional observables, revealing that the EH itself can undergo quantum phase transitions.
 At high frequencies, we find that the EH gains an intrinsic timescale decoupled from the drive, rendering its Floquet-periodic transitions as genuine steady-state features.
-The universal critical behavior persists across different equilibrium phases of the initial state, confirming the transitions’ fundamentally non-equilibrium character, decoupled from the underlying equilibrium phase diagram.
-Figure 1: Temporal entanglement transitions in the driven TFIM.
+The universal critical behavior persists whenever the above Z2-symmetry condition and sector-weight equality are satisfied (including across different equilibrium phases that preserve Z2), confirming a fundamentally non-equilibrium character decoupled from the underlying equilibrium phase diagram (see Supplemental Material (SM) [66] for details).
+Figure 1: Temporal entanglement transitions (TET) in the driven TFIM.
 (a) Schmidt gap $\Delta\lambda=\lambda_{0}-\lambda_{1}$ closing periodically at critical times.
@@
 (c) Subsystem parity expectations showing dynamical spontaneous $\mathbb{Z}_{2}$ symmetry breaking in the entanglement ground state (green) and complementary behavior in the first-excited state (orange).
-The alternating parity pattern provides consistent physical explanation of the echo’s behavior.
-Red vertical lines mark critical times of symmetry breaking, and overlaid across all panels to provide a synchronized, consistent picture.
+The alternating parity pattern physically underlies the echo’s behavior.
+Red vertical lines mark critical times of symmetry breaking, and are overlaid across all panels to provide a synchronized, consistent picture.
 Parameters: $L=24$, $L_{A}=9$, $J=1.0$, $h_{0}=2.0$, $\omega=5.0$, $dt=0.01$.
@@
 where $J$ is the Ising coupling, $h(t)=(h_{0}/2)\cos(\omega t)$ is the oscillating transverse field, and $\sigma_{i}^{\alpha}$ are Pauli matrices.
-We initialize the system in the ground state of the static Hamiltonian ${\cal H}_{\text{static}}={\cal H}(t)|_{t=0}$ and evolve under ${\cal H}(t)$ using time-dependent variational principle (TDVP) [81, 82, 83] for matrix product states (MPS) [84], as implemented in the ITensors library [85] (see Supplemental Material (SM) for details [86]).
-Note that our results are independent of the specific choice of Floquet drive.
+We initialize the system in the ground state of the static Hamiltonian ${\cal H}_{\text{static}}={\cal H}(t)|_{t=0}$ and evolve under ${\cal H}(t)$ using time-dependent variational principle (TDVP) [67, 68, 69] for matrix product states (MPS) [70], as implemented in the ITensors library [71] (see SM [66]).
+Note that our results are independent of the specific Floquet drive.
 Our primary focus is the entanglement dynamics of subsystem $A$, comprising the first $L_{A}$ spins.
@@
 [53]).
-A crossing in the two largest Schmidt values, $\lambda_{0}(t)$ and $\lambda_{1}(t)$, signifies a fundamental reorganization of the ES, corresponding to a non-analyticity in the ground state energy $\epsilon_{0}(t)$ of ${\cal H}_{\text{ent}}(t)$ (with fictitious temperature $T_{\text{fict}}=\lim_{n\to\infty}\frac{1}{n}=0$), signaling a temporal quantum phase transition in the EH.
-Surprisingly, we observe these transitions across all non-zero driving frequencies (from $\omega\ll$ $\min(h_{0},J)$ to $\omega\gg$ $\max(h_{0},J)$), indicating the generic nature of this phenomenon in the periodically-driven TFIM across all frequencies.
-This behavior emerges specifically when the initial state is entangled (such as the ground state of ${\cal H}_{\text{static}}$), rather than a simple product state or a domain wall state with vanishing entanglement at the subsystem boundary (see SM [86]).
+TET are a universal consequence at all driving frequencies provided (i) the driven Hamiltonian and (ii) the initial state preserve the global $\mathbb{Z}_{2}$ symmetry, and (iii) $\rho_{A}(t)$ has support in both subsystem-parity sectors.
+We quantify (iii) by the sector weights $w_{\pm}(t)=\operatorname{Tr}[P_{\pm}\rho_{A}(t)]$, with $P_{\pm}=(\mathds{1}_{A}\pm P_{A})/2$ and $P_{A}=\prod_{i\in A}\sigma_{i}^{x}$.
+Conditions (i)–(ii) imply $\rho_{A}(t),P_{A}]=0$, which we verify using $C(t)=\|[\rho_{A}(t),P_{A}]\|$.
+Within this symmetry-preserving regime, the critical times $t_{c}^{(k)}$ are selected by the operational criterion $w_{+}(t_{c}^{(k)})=w_{-}(t_{c}^{(k)})$ (within the evolution window), at which the Schmidt gap closes and the leading Schmidt vectors exchange subsystem-parity discontinuously.
+By contrast, an isolated suppression of the Schmidt gap or a smooth parity drift—even if accompanied by a sector-weight crossing—indicates a rounded crossover rather than a TET (see Sec.
+III of the SM [66]).
+A crossing in the two largest Schmidt values, $\lambda_{0}(t)$ and $\lambda_{1}(t)$, signifies a fundamental reorganization of the ES, corresponding to a non-analyticity in the ground state energy $\epsilon_{0}(t)$ of ${\cal H}_{\text{ent}}(t)$ (with fictitious temperature $T_{\text{fict}}=\lim_{n\to\infty}\frac{1}{n}=0$), signaling a sharp temporal quantum phase transition in the EH.
+Surprisingly, we observe this behavior from the adiabatic to the high-frequency regime.
+This behavior emerges when the global $\mathbb{Z}_{2}$ symmetry is preserved by the driven Hamiltonian and respected by the initial state, with critical times occurring at sector-weight crossings $w_{+}(t_{c}^{(k)})=w_{-}(t_{c}^{(k)})$.
+Conversely, $\mathbb{Z}_{2}$-breaking initial states (including symmetry-broken ferromagnetic ground states as well as product states such as domain wall and random product states) do not exhibit TET, independent of the drive protocol (see Sec.
+III in SM [66]).
 To diagnose dynamical transitions within the entanglement spectrum, we introduce and track specific observables designed to detect symmetry breaking.
@@
 To directly test this, we compute the expectation values of symmetry operators within the entanglement eigenstates.
-Specifically, we monitor the subsystem parity $\langle\lambda_{n}(t)|P_{A}|\lambda_{n}(t)\rangle$, where $P_{A}=\prod_{i\in A}\sigma_{i}^{x}$.
-A spontaneous change in the parity of the dominant state $|\lambda_{0}(t)\rangle$, coinciding with a vanishing entanglement echo and a vanishing Schmidt gap $\lambda_{0}-\lambda_{1}$, constitutes the hallmark of a temporal entanglement transition.
+Specifically, we monitor the subsystem parity $\langle\lambda_{n}(t)|P_{A}|\lambda_{n}(t)\rangle$.
+A spontaneous change in the parity of the dominant state $|\lambda_{0}(t)\rangle$, coinciding with a vanishing entanglement echo and a vanishing Schmidt gap $\lambda_{0}-\lambda_{1}$, constitutes the hallmark of a TET.
 For each subsequent interval, the dynamics alternate between two dwell times: $T_{-}$, during which the dominant state carries $\langle P_{A}\rangle=-1$, and $T_{+}$, during which it carries $\langle P_{A}\rangle=+1$.
@@
 This synchronization and saturation indicate complete Floquet inheritance: the EH dynamics are governed by a high-frequency effective (Floquet-Magnus) description, in which the entanglement transitions form a uniformly spaced temporal lattice set by an emergent, drive-induced time-scale rather than the bare period.
-See Table I and Sec.
-III.B in SM [86] for further details.
+See Table II, Sec.
+III and Sec.
+IV.B in SM [66] for further details.
 Unless otherwise noted, we use $L=24$ (open boundaries), $L_{A}=4$1212, $J=1.0$, $h_{0}=2.0$ (corresponding to equilibrium criticality), and $\omega=5.0$.
 Time steps $dt\leq 0.1/\omega$ ensure numerical stability; convergence is verified in SM [66].
-Temporal Entanglement Transitions.— The combined signatures of a temporal entanglement transition are unambiguously observed in the dynamics of the driven TFIM.
-Fig.
-1 presents a comprehensive view of the entanglement dynamics for a subsystem of size $L_{A}=9$, revealing periodic occurrences of the transition.
-The onset of a transition is heralded by a critical slowdown of the entanglement spectrum, manifested as a sharp narrowing of the Schmidt gap $\Delta\lambda=\lambda_{0}-\lambda_{1}$ (Fig.
-1a).
-The gap closes to within numerical precision at multiple critical times $t_{c}^{(k)}$}, where we denote the first critical time as $t_{c}^{(k=1)}\equiv t^{*}$, indicating a near-degeneracy of the two largest Schmidt values.
-This closing gap is the direct signature of an impending energy-level crossing in the EH ${\cal H}_{\text{ent}}(t)$ (recall that the largest and the second largest Schmidt values directly correspond to the ground state and first-excited state energies of ${\cal H}_{\text{ent}}(t)$).
-Concurrently, the entanglement echo $E(t)=\left\langle\lambda_{0}(0)\mid\lambda_{0}(t)\right\rangle$ exhibits sharp vanishing at odd critical times (Fig.
-1b).
-At these times $t_{c}^{(k=\text{odd})}$, where the Schmidt gap closes, $|E(t)|^{2}$ drops to zero, establishing that the instantaneous entanglement ground state $|\lambda_{0}(t_{c})\rangle$ orthogonalizes relative to its initial configuration.
-This confirms the reorganization of quantum correlations and the transition into a distinct symmetry sector of the Hilbert space.
-The nature of this sector is revealed by the expectation value of the subsystem parity operator $P_{A}$ (Fig.
-1c).
-For $t<t^{*}$, the ground state of ${\cal H}_{\text{ent}}(t)$ exhibits a well-defined parity of $\langle P_{A}\rangle=+1$.
-At each transition point $t^{*}$, this parity expectation flips discontinuously to $\langle P_{A}\rangle=-1$, indicating a dynamical spontaneous breaking of the $\mathbb{Z}_{2}$ symmetry by the entanglement ground state.
-Notably, the first excited state $|\lambda_{1}(t)\rangle$ displays the opposite parity, confirming that the gap closing corresponds to an exchange of roles between two symmetry-broken states.
-The subsystem magnetization $\langle M_{A}\rangle$ [80] and Loschmidt echo rate function [87, 88, 89, 90] (verified in SM [86]) remain smooth through these events, confirming the transition is of a purely entanglement-driven, dynamical spontaneous symmetry-breaking character.
-We observe these transition signatures across all non-zero driving frequencies $\omega$ studied, indicating the generic nature of this phenomenon in periodically driven Ising chain.
-This behavior emerges specifically when the initial state is entangled (such as the ground state of ${\cal H}_{\text{static}}$), rather than a simple non-entangled state (see SM [86]).
-While transitions occur at all frequencies, the periodic regularity of critical times $t_{c}^{(k)}$} becomes pronounced only at higher driving frequencies, suggesting a crossover to Floquet-like behavior at the level of the EH (details provided in the SM [86]).
-The synchronization of these three diagnostics, namely the closing of the Schmidt gap, the vanishing of the entanglement echo, and the discontinuous flip in parity, provides definitive evidence of a temporal entanglement transition.
+Temporal Entanglement Transitions.— The combined signatures of a TET are unambiguously observed in the driven TFIM (Fig.
+1, $L_{A}=9$).
+The Schmidt gap $\Delta\lambda=\lambda_{0}-\lambda_{1}$ (Fig.
+1a) closes to numerical precision at critical times $t_{c}^{(k)}$ (with first critical time $t_{c}^{(k=1)}\equiv t^{*}$), signaling energy-level crossing in ${\cal H}_{\text{ent}}(t)$.
+Concurrently, the entanglement echo $E(t)=\langle\lambda_{0}(0)\mid\lambda_{0}(t)\rangle$ (Fig.
+1b) vanishes at odd $t_{c}^{(k=\text{odd})}$, confirming that $|\lambda_{0}(t_{c})\rangle$ orthogonalizes into a distinct symmetry sector.
+The subsystem parity $\langle P_{A}\rangle$ (Fig.
+1c) flips discontinuously at each $t_{c}^{(k)}$ between +1+1 and 1-1, evidencing dynamical spontaneous $\mathbb{Z}_{2}$ breaking with respect to the largest Schmidt vector $|\lambda_{0}(t)\rangle$, while the first excited state $|\lambda_{1}(t)\rangle$ carries opposite parity.
+We observe these signatures across all non-zero driving frequencies, indicating the generic nature of this phenomenon in the periodically driven Ising chain.
+While the transitions remain sharp, the spacing of critical times $t_{c}^{(k)}$} becomes nearly periodic only at higher driving frequencies, indicating a crossover in the timing (Floquet inheritance) rather than in the transition itself (see SM [66]).
+The synchronization of these three diagnostics: the closing of the Schmidt gap, the vanishing of the entanglement echo, and the discontinuous flip in parity, provides definitive evidence of a TET.
 These are not independent events but are intrinsically linked manifestations of the same underlying phenomenon: a periodically-driven, dynamical quantum phase transition of the EH ${\cal H}_{\text{ent}}(t)$.
-Finite-Size Scaling.— The temporal entanglement transitions are marked by non-analytic kinks in the ground state energy $\epsilon_{0}(t)=-\ln\lambda_{0}(t)$ of the EH.
-To establish the critical nature and universality of temporal entanglement transitions, we perform comprehensive finite-size scaling analysis across subsystem sizes $L_{A}=4,5,\ldots,12$.
+Finite-Size Scaling.— The TET are marked by non-analytic kinks in the ground state energy $\epsilon_{0}(t)=-\ln\lambda_{0}(t)$ of the EH.
+To establish the critical nature and universality of TET for the critical initial state (where all three conditions are robustly satisfied in the short-time dynamics), we perform comprehensive finite-size scaling analysis across subsystem sizes $L_{A}=4-12$.
 Fig.
@@
 Surprisingly, the correlation length exponent $\nu=1$ establishes this as a continuous quantum phase transition for all ranges of driving frequencies, belonging to the same universality class as the equilibrium 2D classical Ising/1D quantum TFIM.
-Details provided in the SM [86] confirm that this exponent match is coincidental: to the best of our knowledge, temporal entanglement transitions appear to occur independently of the equilibrium phase diagram.
-The universality of these transitions is demonstrated through data collapse using the scaling ansatz (recall $\epsilon_{0}=S_{\text{min}}$):
-$\frac{\epsilon_{0}}{L_{A}}=\frac{1}{L_{A}^{a}}\mathcal{F}\left[\left(\frac{t}{L_{A}}-\frac{t_{c}}{L_{A}}\right)L_{A}^{1/\nu}\right],$ (2)
+Note that this exponent match $\nu=1$ appears coincidental; TET seem to emerge independently of the equilibrium phase diagram, as confirmed in SM [66].
+The universality of these transitions for the critical initial state is demonstrated through data collapse using the scaling ansatz (recall $\epsilon_{0}=S_{\text{min}}$):
+$\frac{\epsilon_{0}}{L_{A}}=\frac{1}{L_{A}^{a}}\mathcal{F}\left[\left(\frac{t}{L_{A}}-\frac{t^{*}}{L_{A}}\right)L_{A}^{1/\nu}\right],$ (2)
 where $\mathcal{F}$ is a universal scaling function, and $\nu$ is the critical exponent corresponding to diverging correlation length.
+For our critical initial state, $t^{*}$ occurs within the simulated time window for all driving frequencies studied, enabling a uniform finite-size scaling analysis from the adiabatic to the high-frequency regime.
 Fig.
@@
 This universality across multiple transitions suggests a common underlying fixed point.
-The exponent $\nu\simeq 1.00$ defines a new universality class for non-equilibrium entanglement dynamics in driven systems, which appears to be completely decoupled from the underlying equilibrium criticality (more on this below).
-Figure 2: Finite-size scaling analysis of temporal entanglement transitions.
+For the critical initial state, the exponent $\nu\simeq 1.0$ defines a new universality class for non-equilibrium entanglement dynamics in driven systems, which appears to be completely decoupled from the underlying equilibrium criticality (more on this below).
+Figure 2: Finite-size scaling analysis of TET starting with critical initial state $(h_{0}/2=1.0)$.
 (a) Critical time scaling $t^{*}/L_{A}\propto L_{A}^{-1/\nu}$ with $\nu=1.00$.
@@
 Parameters: $L=24$, $J=1.0$, $h_{0}=2.0$, $\omega=5.0$, $dt=0.01$.
-Figure 3: Frequency dependence of the first critical time $t^{*}$ across subsystem sizes $L_{A}=4-12$.
+Figure 3: Frequency dependence of the first critical time $t^{*}$ across subsystem sizes $L_{A}=4-12$ for critical initial state $(h_{0}/2=1.0)$.
 At low frequencies, $t^{*}\propto\omega^{-1}$ (adiabatic regime), while at high frequencies ($\omega\gtrsim 10$), $t^{*}$ saturates to frequency-independent values determined solely by subsystem size, indicating a crossover to Floquet steady-state behavior where the EH develops an intrinsic timescale.
-The universal saturation indicates that temporal entanglement transitions become intrinsic properties of the effective time-independent dynamics rather than driven phenomena, validating the Floquet-Magnus description and establishing characteristic timescales independent of the external driving protocol.
+The universal saturation indicates that TET become intrinsic properties of the effective time-independent dynamics rather than driven phenomena, validating the Floquet-Magnus description and establishing characteristic timescales independent of the external driving protocol.
 Parameters: $L=24$, $J=1.0$, $h_{0}=2.0$.
@@
 9 of SM [66]).
-Effective Steady State at High Frequency.— At high driving frequencies, the temporal entanglement transitions can be understood through Floquet-Magnus theory [91, 92, 93].
+Effective Steady State at High Frequency.— At high driving frequencies, TET acquire an intrinsic timescale explained by Floquet-Magnus theory [72, 73, 74].
 We decompose the time-dependent Hamiltonian as $\mathcal{H}(t)=\mathcal{A}+\mathcal{B}\cos(\omega t)$, where $\mathcal{A}=-J\sum_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}$ and $\mathcal{B}=-(h_{0}/2)\sum_{i}\sigma_{i}^{x}$.
@@
 Beyond renormalizing the Ising coupling, the expansion generates new terms: YY interactions, enhanced transverse fields at boundaries, and crucially, three-body interactions $\sim\sigma_{i-1}^{z}\sigma_{i}^{x}\sigma_{i+1}^{z}$ that couple neighboring bonds through local spin flips.
-As demonstrated in SM [86], exact time evolution under $\mathcal{H}(t)$ shows good agreement with evolution under $\mathcal{H}_{\text{eff}}$ at higher frequencies, with entanglement entropy, Schmidt gap dynamics, and state fidelity exhibiting near-perfect overlap.
-Temporal entanglement transitions persist in the effective evolution with identical critical exponent $\nu$, confirming that these phenomena represent genuine features of the driven steady state where the EH develops an intrinsic timescale.
-Table 1: Conditions for temporal entanglement transitions across different scenarios.
-System
-Initial state
-(Product/Entangled)
-Transitions
-Equilibrium TFIM Any No
-Driven TFIM Product No
-Driven TFIM Entangleda Yes
-Conserved‐charge
-models
-Any Yesb
-a See SM for detailed discussion on dependence on initial state. b Only when $\rho_{A}$ is forced between disconnected sectors [51, 52].
-Discussion and Conclusion.— We have demonstrated that the Floquet driven Ising chain hosts temporal entanglement transitions characterized by dynamical spontaneous symmetry breaking within the EH.
-These transitions exhibit universal critical behavior with exponent $\nu=1$, identical to equilibrium quantum criticality in the transverse-field Ising universality class, yet emerge from non-equilibrium dynamics that are decoupled from static criticality, as shown by their persistence across different equilibrium phases ($h_{0}=1.6$ to $2.4$, see SM [86]).
-This work establishes that: (1) the ES develops non-analyticities for all $\omega$, with periodic reorganizations at higher frequencies; (2) transitions are diagnosed by synchronized Schmidt gap closure, entanglement echo vanishing, and parity flipping, evidencing dynamical spontaneous symmetry breaking; (3) finite-size scaling reveals universal $\nu=1$ critical behavior across frequencies; (4) at high frequencies, Floquet-Magnus theory provides an effective description accurately capturing the transition; (5) as summarized in Table 1, transitions require both driving and initial entanglement, distinguishing them from transitions found for initial product states [53] (more details in the SM [86]).
+As demonstrated in SM [66], exact time evolution under $\mathcal{H}(t)$ shows good agreement with that under $\mathcal{H}_{\text{eff}}$ at higher frequencies, with entanglement entropy, Schmidt gap dynamics, and state fidelity exhibiting near-perfect overlap.
+TET persist in the effective evolution with identical critical exponent $\nu$, confirming that these phenomena represent genuine features of the driven steady state where the EH develops an intrinsic timescale.
+Table 1: Conditions for TET.
+System Drive $\mathbb{Z}_{2}$ Initial state $\mathbb{Z}_{2}$ TET
+Equilibrium TFIM No
+Driven TFIM Yes No (symmetry-broken) No
+Driven TFIM Yes Yes Yesa
+Conserved-charge models Any Yesb
+a Provided $\rho_{A}(t)$ carries weight in both subsystem-parity sectors, with TET occurring when $w_{+}(t_{c}^{(k)})=w_{-}(t_{c}^{(k)})$ while $\rho_{A},P_{A}]$ remains negligible, see SM Sec.
+III.
+Many standard product-state preparations are $\mathbb{Z}_{2}$-breaking and thus excluded, while $\mathbb{Z}_{2}$-symmetric (often entangled) initial states are allowed.
+b Only when $\rho_{A}$ is dynamically forced between disconnected sectors [52, 53].
+Discussion and Conclusion.— We have demonstrated that the Floquet driven Ising chain hosts temporal entanglement transitions (TET) characterized by dynamical spontaneous symmetry breaking within the EH.
+These transitions exhibit universal critical behavior with exponent $\nu=1$ for critical initial state ($h_{0}/2=1.0$), identical to equilibrium quantum criticality in the transverse-field Ising universality class, yet emerge from non-equilibrium dynamics that are decoupled from static criticality, as shown by their persistence across different equilibrium phases ($h_{0}=1.6$ to $2.4$, see SM [66]).
+This work establishes that: (1) the ES develops non-analyticities for all $\omega$, with periodic reorganizations at higher frequencies; (2) transitions are diagnosed by synchronized Schmidt gap closure, entanglement echo vanishing, and parity flipping, evidencing dynamical spontaneous symmetry breaking; (3) finite-size scaling for the critical initial state reveals universal $\nu=1$ critical behavior across frequencies; (4) at high frequencies, Floquet-Magnus theory provides an effective description accurately capturing the transition; (5) TET occur precisely when the sector weights become equal, $w_{+}(t_{c}^{(k)})=w_{-}(t_{c}^{(k)})$, provided the driven Hamiltonian and initial state preserve global $\mathbb{Z}_{2}$ symmetry (Table 1; also see SM, Sec.
+III [66]).
+Experimental access to these transitions is feasible: the key observables (Schmidt gap, parity expectations, entanglement echo) require only low-lying eigenstates of the EH, measurable via randomized measurements [75, 76] or Ramsey spectroscopy [77].
 Fig.
@@
 The collapse of $t^{*}$ curves across subsystem sizes demonstrates universal Floquet inheritance, where the EH acquires periodicity and full data collapse is satisfied by the same $\nu=1$ exponent for all transitions [66].
-These findings establish temporal entanglement transitions as fundamental aspects of non-equilibrium Floquet quantum dynamics.
+These findings establish TET as fundamental aspects of non-equilibrium Floquet quantum dynamics.
 Notably, previous work on Floquet symmetry-protected topological system has demonstrated band crossings in entanglement spectrum [78], however the question of robust universality has been an open question.
-This is what we have addressed in this work, where our temporal entanglement transitions arise universally for any initially entangled state (topological or trivial) and exhibit nonanalytic reorganization (quantum phase transition) of the EH with Ising‐class critical exponents $\nu=1$.
-The observed behavior suggests driven systems can host rich critical phenomena as diagnosed through their entanglement dynamics.
-Temporal entanglement transitions define a class of non-equilibrium criticality distinct from Loschmidt echo singularities and magnetization dynamics (see SM [86]).
+This is what we have addressed in this work, where our TET arise whenever the driven Hamiltonian and initial state respect global $\mathbb{Z}_{2}$, with critical times determined by sector-weight crossings $w_{+}(t_{c}^{(k)})=w_{-}(t_{c}^{(k)})$, yielding a nonanalytic reorganization (quantum phase transition) of the EH with Ising-class critical exponents $\nu=1$.
+TET define a class of non-equilibrium criticality distinct from Loschmidt echo singularities and magnetization dynamics (see SM [66]).
 The exact Jordan-Wigner equivalence between the TFIM and Kitaev chain suggests that these transitions should also manifest in driven topological superconductors, opening pathways to entanglement-based probes of Floquet Majorana physics [79, 80].
-The requirement for initial entangled states motivates future exploration of how entanglement resources control critical structures in quantum resource theories [97].
-Future directions include extensions to higher dimensions, long-range interactions, and potential realization in cold-atoms or trapped ion experiments [98, 99, 100].
+The requirement of symmetry-preserving initial conditions with critical times selected by sector-weight degeneracies motivates future exploration of how symmetry and entanglement resources jointly control non-equilibrium criticality structures in quantum resource theories [81].
+Beyond the TFIM, the criterion (global symmetry, symmetric initial state, sector-weight equality $w_{+}(t)=w_{-}(t)$) suggests analogous transitions in other driven systems with discrete symmetries, including higher-spin chains, symmetry-protected topological phases, and lattice gauge theories, with experimental prospects in cold-atom and trapped-ion platforms [82, 83, 84, 75].
 Acknowledgments.— K.
@@
 493420525, via large equipment grants (GOEGrid).
-This work is supported by the Office of Advanced Scientific Computing Research, Exploratory Research for Extreme Scale Science (EXPRESS) program, Office of Science of the U.S.
-Department of Energy under Award Number DE-SC0026216 (A.
-P.).
+This material is based upon work supported by the U.S.
+Department of Energy, Office of Science, Office of Advanced Scientific Computing Research via the Exploratory Research for Extreme Scale Science (EXPRESS) program under Award Number DE-SC0026216.
 This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
@@
 P.
-thanks the Kavli Institute for Theoretical Physics (KITP) for its hospitality during the program on “Noise-robust Phases of Quantum Matter,” during which part of this work was completed.
+thanks the Kavli Institute for Theoretical Physics (KITP) for its hospitality during the program on “Noise-robust Phases of Quantum Matter”, during which part of this work was completed.
 Data and code availability.— All data and code used for data generation are available on Zenodo on reasonable request [85].
+Disclaimer.— This report was prepared as an account of work sponsored by an agency of the United States Government.
+Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights.
+Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof.
+The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.

### v3

@@
 At high frequencies, we find that the EH gains an intrinsic timescale decoupled from the drive, rendering its Floquet-periodic transitions as genuine steady-state features.
-The universal critical behavior persists whenever the above Z2-symmetry condition and sector-weight equality are satisfied (including across different equilibrium phases that preserve Z2), confirming a fundamentally non-equilibrium character decoupled from the underlying equilibrium phase diagram (see Supplemental Material (SM) [66] for details).
+The universal critical behavior persists whenever the above $\mathbb{Z}_{2}$-symmetry condition and sector-weight equality are satisfied (including across different equilibrium phases that preserve $\mathbb{Z}_{2}$), confirming a fundamentally non-equilibrium character decoupled from the underlying equilibrium phase diagram (see Supplemental Material (SM) [1] for details).
 Figure 1: Temporal entanglement transitions (TET) in the driven TFIM.
@@
 The requirement of symmetry-preserving initial conditions with critical times selected by sector-weight degeneracies motivates future exploration of how symmetry and entanglement resources jointly control non-equilibrium criticality structures in quantum resource theories [18].
-Beyond the TFIM, the criterion (global symmetry, symmetric initial state, sector-weight equality $w_{+}(t)=w_{-}(t)$) suggests analogous transitions in other driven systems with discrete symmetries, including higher-spin chains, symmetry-protected topological phases, and lattice gauge theories, with experimental prospects in cold-atom and trapped-ion platforms [82, 83, 84, 75].
+Beyond the TFIM (despite its free-fermion Jordan-Wigner mapping), our symmetry/sector-competition criterion (global symmetry, symmetric initial state, and sector-weight equality $w_{+}(t)=w_{-}(t)$) suggests analogous transitions in interacting $\mathbb{Z}_{2}$-symmetric chains and, more broadly, in other driven systems with discrete symmetries (higher-spin chains, symmetry-protected topological phases, and lattice gauge theories), with experimental prospects in cold-atom and trapped-ion platforms [25, 9, 42, 11].
 Acknowledgments.— K.